# On Championing Foundation Models: From Explainability to Interpretability

### v1

@@ 3.1.1 Interpreting Through Generalization Analysis
The term $\mathcal{N}(\mathcal{A},\rho,\epsilon)$ denotes the $\epsilon$-covering number of the hypothesis set $\mathcal{A}$, utilized to regulate model complexity. In this context, $\rho$ represents the prediction disparity between the two algorithms on the most adverse data sequence, governed by the Lipschitz constant of the transformer design. The bound in 3.6 attains a rate of $1/\sqrt{nT}$ by encompassing the algorithm space with a resolution of $\varepsilon$. For Lipschitz architectures possessing $\operatorname{dim}(\mathcal{A})$ trainable weights, it follows that $\log\mathcal{N}(\mathcal{A},\rho,\varepsilon)\sim\operatorname{dim}(\mathcal{H})\log(1/\varepsilon)$. Consequently, the excess risk is constrained by $\sqrt{\frac{\operatorname{dim}(\mathcal{A})}{nT}}$ up to logarithmic factors, and will diminish as $n$ and $T$ go toward infinity.
@@ 3.1.2 Interpreting Through Expressivity Analysis
(Linking Gradient Descent and Self-Attention [103]) Given a single-head linear attention layer and a set of in-context example ${e}_{j}=\left({{x}_{j},{y}_{j}}\right)$ for $j=1,\ldots,N$, it is possible to construct key, query, and value matrices ${W}_{K},{W}_{Q},{W}_{V}$, as well as a projection matrix $P$, such that a transformer step applied to each in-context example ${e}_{j}$ is equivalent to the gradient-induced update ${e}_{j}\leftarrow\left({{x}_{j},{y}_{j}}\right)+\left({0,-{\Delta W}{x}_{j}}\right)=\left({{x}_{j},{y}_{j}}\right)+{PV}{K}^{T}{q}_{j}$, resulting in ${e}_{j}=\left({{x}_{j},{y}_{j}-\Delta{y}_{j}}\right)$. This dynamic is similarly applicable to the query token $\left({{x}_{N+1},{y}_{N+1}}\right)$
@@
The studies [96] and [91] conducted further investigation into the implementation of ICL for linear regression, specifically through the utilization of single-layer linear attention models. They demonstrated that a global minimizer of the population ICL loss can be interpreted as a single gradient descent step with a matrix step size. [106] provided a similar result specifically for ICL in isotropic linear regression. It is noteworthy that [91] also addressed the optimization of attention models; however, their findings necessitate an infinite number of pretraining tasks. Similarly, [105] established task complexity bounds for pretraining based on uniform convergence and demonstrated that pretrained transformers are capable of performing algorithm selection. Furthermore, a simplified linear parameterization was utilized in [110], which provided more refined task complexity boundaries for the purpose of pretraining a single-layer linear attention model.
That said, there remain areas in existing research that require refinement. For example, many current studies attempt to model each gradient descent step using a single attention layer within transformer architectures. This approach raises questions regarding its suitability, particularly given that transformers are typically not extremely deep. For instance, variants of models like GPT-2 and BERT usually consist of 12 to 48 layers, which constrains the number of gradient descent steps that can be effectively modeled. This limitation suggests that further investigation is needed to refine these models and better align them with the theoretical underpinnings of gradient descent in the context of ICL.
@@ 3.1.3 Concluding Remarks
Constructing Effective In-Context Examples: Another major challenge for users is how to construct appropriate in-context examples as prompts, as shown in Figure 5. Numerous theoretical studies have provided insights into this challenge. Research [93, 92] suggests that longer prompts generally lead to lower generalization errors and improved performance because they offer richer contextual information that aids the model in better understanding the task. Additionally, [110] found that model performance improves when the context length during inference closely matches that of pretraining. Furthermore, when in-context examples share similar patterns or structures with the query, the model’s performance can be significantly enhanced. Specifically, the amount of influence that various parameters have on the performance of ICL generalization was quantified by [92]. These factors include the quantity of relevant features and the fraction of context instances that include the same relevant pattern as the new query.
@@ 3.2.1 Interpreting Through Generalization Analysis
${\mathbf{E}}_{i}=\left(\begin{matrix}{\mathbf{x}}_{i}&{\mathbf{y}}_{i,1}&\cdots&{\mathbf{y}}_{i,K-1}\\ {\mathbf{y}}_{i,1}&{\mathbf{y}}_{i,2}&\cdots&{\mathbf{y}}_{i,K}\end{matrix}\right),{\mathbf{Q}}_{k}=\left(\begin{matrix}{\mathbf{z}}_{0}&{\mathbf{z}}_{1}&\cdots&{\mathbf{z}}_{k-2}&{\mathbf{z}}_{k-1}\\ {\mathbf{z}}_{1}&{\mathbf{z}}_{2}&\cdots&{\mathbf{z}}_{k-1}&\mathbf{0}\end{matrix}\right),i\in\left[{n}_{tr}\right],$
@@
where $n$ represents the number of prompt-label pairs $\left\{\mathbf{P}^{i},\mathbf{z}^{i}\right\}_{i=1}^{n}$ and $\boldsymbol{\theta}$ denotes model weights. The loss function utilized is the squared loss, specifically given by $\ell\left(\mathbf{P}^{i},\mathbf{z}^{i};\boldsymbol{\theta}\right)=\frac{1}{2}\left\|\mathbf{z}^{i}-F\left(\mathbf{P}^{i};\boldsymbol{\theta}\right)\right\|^{2}$ where $F\left(\mathbf{P}^{i};\boldsymbol{\theta}\right)$ denotes the output of the transformer model. Furthermore, following the works of [99, 97, 98, 101], the CoT generalization error for a testing query $\mathbf{x}_{\text{query }}$, the testing data distribution $\mathcal{D}^{\prime}$, and the labels $\left\{\mathbf{z}_{k}\right\}_{k=1}^{K}$ on a $K$-step testing task $f\in\mathcal{T}^{\prime}$ is defined as
$R_{CoT,\mathbf{x}_{\text{query }}\sim\mathcal{D}^{\prime},f\in\mathcal{T}^{\prime}}^{f}(\boldsymbol{\theta})=\mathbb{E}_{\mathbf{x}_{\text{query }}\sim\mathcal{D}^{\prime}}\left[\frac{1}{K}\sum_{k=1}^{K}1\left[\mathbf{z}_{k}\neq\mathbf{v}_{k}\right]\right]$
@@
${n}\geq\Omega\left({\alpha}^{-1}\right)$
@@
$\mu_{j}^{\prime}\in\operatorname{span}\left(\mu_{1},\mu_{2},\cdots,\mu_{M}\right)$
@@
$n_{ts}^{f}\geq\Omega\left(\left(\alpha^{\prime}\tau^{f}\rho^{f}\right)^{-2}\log M\right)$ (5)
@@ 3.3.1 Interpreting Through Generalization Analysis
Additionally, [94] propose a PAC-based framework for in-context learnability and utilize it to present the finite sample complexity results. They utilize their framework to demonstrate that, under modest assumptions, when the pretraining distribution consists of a combination of latent tasks, these tasks can be effectively acquired by ICL. This holds even when the input significantly diverges from the pretraining distribution. In their demonstration, they show that the ratio of the prompt probabilities, which is dictated by the ground-truth components, converges to zero at an exponential rate. This rate depends on the quantity of in-context examples as well as the minimum Kullback–Leibler divergence between the ground-truth component and other mixture components, indicating the distribution drift.
@@ 4.1.2 Concluding Remarks
However, the current understanding of emergent abilities is still in its infancy, particularly in terms of rigorous theoretical analysis of scaling laws. This reveals a gap in our knowledge of the relationship between model size and emergent capabilities, which warrants further exploration. Despite limited existing research, we posit that a promising avenue for addressing this challenge lies in mechanistic explanations grounded in mathematical analysis or control theory, such as the exploration of phase transitions in complex systems, and lévy process driven Stochastic Differential Equation (SDEs).
@@ 4.2.1 Interpretation for Pretraining Dynamics
Transformer architectures have excelled across various research domains, becoming the backbone of many FMs. A critical aspect of interpreting these models lies in understanding their training dynamics. Recent research has focused on how representations emerge from gradient-based optimization, particularly emphasizing the dynamics of individual attention layers—the core components of transformers.
@@
Expanding on the theoretical underpinnings of these mechanisms, [85, 86] established a compelling equivalence between the optimization dynamics of a singular attention layer and a particular Support Vector Machine (SVM) issue. This relationship offered a novel perspective on attention layers, indicating that they may utilize analogous optimization concepts to those present in conventional machine learning algorithms, so bridging the gap between classical techniques and modern deep learning frameworks.
@@ 4.3.2 Concluding Remarks
In this chapter, we primarily utilize theoretical results to interpret the underlying mechanisms of RLHF, thereby addressing the challenges. Additionally, we explore potential directions for future theoretical advancements.

### v2

@@ 4.1.1 Interpreting Through Generalization Analysis
The term $\mathcal{N}(\mathcal{A},\rho,\epsilon)$ denotes the $\epsilon$-covering number of the hypothesis set $\mathcal{A}$, utilized to regulate model complexity. In this context, $\rho$ represents the prediction disparity between the two algorithms on the most adverse data sequence, governed by the Lipschitz constant of the transformer design. The bound in 4.6 attains a rate of $1/\sqrt{nT}$ by encompassing the algorithm space with a resolution of $\varepsilon$. For Lipschitz architectures possessing $\operatorname{dim}(\mathcal{A})$ trainable weights, it follows that $\log\mathcal{N}(\mathcal{A},\rho,\varepsilon)\sim\operatorname{dim}(\mathcal{H})\log(1/\varepsilon)$. Consequently, the excess risk is constrained by $\sqrt{\frac{\operatorname{dim}(\mathcal{A})}{nT}}$ up to logarithmic factors, and will diminish as $n$ and $T$ go toward infinity.
@@ 4.1.2 Interpreting Through Expressivity Analysis
(Linking Gradient Descent and Self-Attention [97]) Given a single-head linear attention layer and a set of in-context example ${e}_{j}=\left({{x}_{j},{y}_{j}}\right)$ for $j=1,\ldots,N$, it is possible to construct key, query, and value matrices ${W}_{K},{W}_{Q},{W}_{V}$, as well as a projection matrix $P$, such that a transformer step applied to each in-context example ${e}_{j}$ is equivalent to the gradient-induced update
${e}_{j}\leftarrow\left({{x}_{j},{y}_{j}}\right)+\left({0,-{\Delta W}{x}_{j}}\right)=\left({{x}_{j},{y}_{j}}\right)+{PV}{K}^{T}{q}_{j},$
resulting in ${e}_{j}=\left({{x}_{j},{y}_{j}-\Delta{y}_{j}}\right)$. This dynamic is similarly applicable to the query token $\left({{x}_{N+1},{y}_{N+1}}\right)$
@@
[101] and [102] conducted further investigation into the implementation of ICL for linear regression, specifically through the utilization of single-layer linear attention models. They demonstrated that a global minimizer of the population ICL loss can be interpreted as a single gradient descent step with a matrix step size. [103] provided a similar result specifically for ICL in isotropic linear regression. It is noteworthy that [102] also addressed the optimization of attention models; however, their findings necessitate an infinite number of pretraining tasks. Similarly, [91] established task complexity bounds for pretraining based on uniform convergence and demonstrated that pretrained transformers are capable of performing algorithm selection. Furthermore, a simplified linear parameterization was utilized in [90], which provided more refined task complexity boundaries for the purpose of pretraining a single-layer linear attention model.
Notably, exploring the connection between gradient descent and ICL is highly valuable for advancing the theoretical understanding of ICL. This is because it allows us to leverage the extensive tools from traditional learning theory, such as the convergence proof frameworks for gradient descent and stochastic gradient descent, as outlined in Theorems 3.10 and 3.11, respectively.
Furthermore, there remain areas in existing research that require refinement. For example, many current studies attempt to model each gradient descent step using a single attention layer within transformer architectures. This approach raises questions regarding its suitability, particularly given that transformers are typically not extremely deep. For instance, variants of models like GPT-2 and BERT usually consist of 12 to 48 layers, which constrains the number of gradient descent steps that can be effectively modeled. This limitation suggests that further investigation is needed to refine these models and better align them with the theoretical underpinnings of gradient descent in the context of ICL.
@@ 4.1.3 Concluding Remarks
Constructing Effective In-Context Examples: Another major challenge for users is how to construct appropriate in-context examples as prompts, as shown in Figure 5. Numerous theoretical studies have provided insights into this challenge. [86, 88] suggests that longer prompts generally lead to lower generalization errors and improved performance because they offer richer contextual information that aids the model in better understanding the task. Additionally, [90] found that model performance improves when the context length during inference closely matches that of pretraining. Furthermore, when in-context examples share similar patterns or structures with the query, the model’s performance can be significantly enhanced. Specifically, the amount of influence that various parameters have on the performance of ICL generalization was quantified by [88]. These factors include the quantity of relevant features and the fraction of context instances that include the same relevant pattern as the new query.
@@ 4.2.1 Interpreting Through Generalization Analysis
and
${\mathbf{E}}_{i}=\left(\begin{matrix}{\mathbf{x}}_{i}&{\mathbf{y}}_{i,1}&\cdots&{\mathbf{y}}_{i,K-1}\\ {\mathbf{y}}_{i,1}&{\mathbf{y}}_{i,2}&\cdots&{\mathbf{y}}_{i,K}\end{matrix}\right),{\mathbf{Q}}_{k}=\left(\begin{matrix}{\mathbf{z}}_{0}&{\mathbf{z}}_{1}&\cdots&{\mathbf{z}}_{k-2}&{\mathbf{z}}_{k-1}\\ {\mathbf{z}}_{1}&{\mathbf{z}}_{2}&\cdots&{\mathbf{z}}_{k-1}&\mathbf{0}\end{matrix}\right),$
@@
where $n$ represents the number of prompt-label pairs $\left\{\mathbf{P}^{i},\mathbf{z}^{i}\right\}_{i=1}^{n}$ and $\boldsymbol{\theta}$ denotes model weights. The loss function utilized is the squared loss, specifically given by
$\ell\left(\mathbf{P}^{i},\mathbf{z}^{i};\boldsymbol{\theta}\right)=\frac{1}{2}\left\|\mathbf{z}^{i}-F\left(\mathbf{P}^{i};\boldsymbol{\theta}\right)\right\|^{2},$
where $F\left(\mathbf{P}^{i};\boldsymbol{\theta}\right)$ denotes the output of the transformer model. Furthermore, following the works [104, 105, 106, 88], the CoT generalization error for a testing query $\mathbf{x}_{\text{query }}$, the testing data distribution $\mathcal{D}^{\prime}$, and the labels $\left\{\mathbf{z}_{k}\right\}_{k=1}^{K}$ on a $K$-step testing task $f\in\mathcal{T}^{\prime}$ is defined as
$R_{CoT,\mathbf{x}_{\text{query }}\sim\mathcal{D}^{\prime},f\in\mathcal{T}^{\prime}}^{f}(\boldsymbol{\theta})=\mathbb{E}_{\mathbf{x}_{\text{query }}\sim\mathcal{D}^{\prime}}\left[\frac{1}{K}\sum_{k=1}^{K}1\left[\mathbf{z}_{k}\neq\mathbf{v}_{k}\right]\right].$
@@
${n}\geq\Omega\left({\alpha}^{-1}\right),$
@@
$\mu_{j}^{\prime}\in\operatorname{span}\left(\mu_{1},\mu_{2},\cdots,\mu_{M}\right),$
@@
$n_{ts}^{f}\geq\Omega\left(\left(\alpha^{\prime}\tau^{f}\rho^{f}\right)^{-2}\log M\right).$ (6)
@@ 4.3.1 Interpreting Through Generalization Analysis
Building on the methodologies outlined in [110, 102], we focus on the design of training prompts structured as
$\left({{x}_{1},h\left({x}_{1}\right),\ldots,{x}_{N},h\left({x}_{N}\right),{x}_{\text{query }}}\right).$
@@
Additionally, [89] adapt the PAC-based framework to analyze in-context learnability and leverage it to derive finite sample complexity results. They utilize their framework to demonstrate that, under modest assumptions, when the pretraining distribution consists of a combination of latent tasks, these tasks can be effectively acquired by ICL. This holds even when the input significantly diverges from the pretraining distribution. In their demonstration, they show that the ratio of the prompt probabilities, which is dictated by the ground-truth components, converges to zero at an exponential rate. This rate depends on the quantity of in-context examples as well as the minimum Kullback–Leibler divergence between the ground-truth component and other mixture components, indicating the distribution drift.
@@ 4.3.2 Concluding Remarks
4.3.3 Future Directions for Theoretical Foundations in Prompt Engineering for FMs
In the preceding section, we highlighted the pivotal role of prompt engineering in expanding the reasoning capabilities of FMs. Foundational advancements, such as ICL and CoT prompting, allow FMs to engage in complex reasoning by guiding them through structured, sequential logic.
Building on these core techniques, an array of prompt engineering strategies has emerged, extending FMs’ capacity to handle increasingly intricate tasks [111]. For example, CoT prompting, introduced by [5], enables FMs to decompose multi-step problems into logical sub-steps, effectively simulating human problem-solving approaches. More recent methods, such as Automatic Chain of Thought (Auto-CoT) prompting [112], automate the creation of reasoning chains, enabling more streamlined and systematic prompt generation through structured cues like “Let’s think step-by-step,” which supports systematic reasoning in complex problem-solving scenarios.
Further innovations include self-consistency, which enhances inference reliability by generating diverse reasoning paths and selecting the most coherent answer, a technique particularly suited to tasks where multiple solution paths are valid. Expanding on this, frameworks such as Tree of Thoughts (ToT) [113] and Graph of Thoughts (GoT) [114] introduce tree-based and graph-based reasoning structures, allowing FMs to explore complex problem spaces with flexibility, backtracking, and exploration—abilities essential for addressing highly complex tasks.
Despite these practical advancements, a comprehensive theoretical foundation explaining the inner workings of prompt engineering remains an open frontier. Future research should prioritize developing formal frameworks that elucidate the mechanisms by which prompt engineering influences FM behavior and inference. Initial theoretical studies, such as those examining retrieval-augmented generation (RAG) [115, 116, 117, 118], have begun to quantify the trade-offs between the benefits of external information and the risks introduced by noise. By modeling RAG as a fusion of the FM’s internal knowledge distribution with the distribution of retrieved texts, these studies offer a theoretical framework for understanding the balance between beneficial knowledge and the risk of misinformation, based on distributional differences in token prediction. While promising, these findings represent only an initial step. Comprehensive theoretical frameworks that can explain a broader range of prompting strategies are still urgently needed.
Establishing deeper theoretical insights into prompt engineering would not only clarify how prompts shape FM reasoning processes but also support the development of more interpretable, reliable FMs for complex real-world applications. Such insights would enable us to harness the full potential of FMs with greater predictability and control, paving the way for robust applications across diverse domains.
@@ 5.1.2 Concluding Remarks
However, the current understanding of emergent abilities is still in its infancy, particularly in terms of rigorous theoretical analysis of scaling laws. This reveals a gap in our knowledge of the relationship between model size and emergent capabilities, which warrants further exploration. Despite limited existing research, we posit that a promising avenue for addressing this challenge lies in mechanistic explanations grounded in mathematical analysis or control theory, such as the exploration of phase transitions from the perspective of dynamic systems.
@@ 5.2.1 Interpretation for Pretraining Dynamics
Transformer architectures have excelled across various research domains, becoming the backbone of many FMs. A critical aspect of interpreting these models lies in understanding their training dynamics. Recent research has focused on how representations emerge from gradient-based optimization, particularly emphasizing the dynamics of individual attention layers—the core components of transformers. Notably, many techniques for theoretically analyzing the gradient-based optimization process have evolved from the proof frameworks for gradient descent and stochastic gradient descent detailed in Theorems 3.10 and 3.11.
@@
Expanding on the theoretical underpinnings of these mechanisms, [124] and [125] established a compelling equivalence between the optimization dynamics of a singular attention layer and a particular Support Vector Machine (SVM) issue. This relationship offered a novel perspective on attention layers, indicating that they may utilize analogous optimization concepts to those present in conventional machine learning algorithms, so bridging the gap between classical techniques and modern deep learning frameworks.
@@ 5.3.2 Concluding Remarks
In this section, we primarily utilize theoretical results to interpret the underlying mechanisms of RLHF, thereby addressing the challenges. Additionally, we explore potential directions for future theoretical advancements.